Consider two projects. The first project pays benefits of $90 today and nothing else. The second project pays nothing today, nothing one year from now, but $100 two years from now. Which project would be preferred if the discount rate were 0%? What if the rate increased to 10%?


1
Expert's answer
2020-09-30T15:04:14-0400

At a discount rate of 0%,

PV1 =90/(1+0)2=$90

the present value calculations would be $90 for project one

Present value of project two is;

PV2=100/(1+0)2= $100

therefore, project two is preferred.

At a rate of 10%, the PV of project one is still $90. For project two,

PV=100/(1.1)2=82.6446

PV is now 82.6; therefore, project one is preferred.